A marksman has 70% accuracy hitting targets at 1000 yards. Find the probability that he will make Exactly 3 of his next 5 shots , exactly 15 of his next 20 shots , at least 5 of his next 6 shots , and no more than 4 of his next 6 shots .
Answer by ewatrrr(24785)

p(hit) = .70
Using a TI calculator 0r similarly a Casio fx-115 ES plus
.....
n = 5, P(x = 3)= 3C5(.70^3)(.30)^2 0r binompdf(5, .70, 3)= .3087 0r 30.87%
....
n = 20, P(x=15)= 20C15(.70^15)(.30)^5 0r binompdf(20, .70, 15)= .1789
.....
n = 6
P(x ≥ 5) = 1 - P(x ≤ 4) = 1 - binomcdf(6,.70,4) = 1- .5798 = .4212
P(x ≤ 4) = binomcdf(6,.70,4)= .5798